# Aritro Pathak

### v1

@@ Abstract
Given any integer $n\geq 3$, $C_{1}>1$, and $0<\beta<1$, there is a constant $\kappa\in(0,1)$, so that the following holds. Define, $N_{\beta}=\lceil\frac{1}{\beta}\rceil$ and $\beta_{1}:=\frac{1}{4(N_{\beta}+1)}$.
@@
Note that, $\beta_{1}\leq\beta/2$. In particular, we will only employ the Eq. 1 condition for any point $x\in\mathbb{R}^{n}$, particularly for planes passing through $x$. We do not restrict this condition to only points $x\in\partial\Omega$. Of course, if we were to choose the point $x$ and the plane $V\ni x$ far from the boundary $\partial\Omega$ depending on $\beta$, then the condition is vacuously true, and so the statement is only interesting when the point $x$ and the chosen plane $V$ are close to the boundary.
@@ 3 Preliminaries
$\sup_{y\in C(x,r)\cap\partial\Omega}\frac{\operatorname{dist}(y,V)}{r}\geq\sup_{y\in B(x,r)\cap\partial\Omega}\frac{\operatorname{dist}(y,V)}{r},\sup_{y\in C(x,\mathscr{r})\cap V}\frac{\operatorname{dist}(y,\partial\Omega)}{r}\geq\sup_{y\in B(x,r)\cap V}\frac{\operatorname{dist}(y,\partial\Omega)}{r}$ (10)
@@ 4 Notation
Throughout the proof, $C_{i}>0$ for any $i\geq 1$, denotes a constant whose value may change from line to line. When necessary, we indicate the dependence of $C$ on ambient parameters. We use multiple different constants which change from line to line depending on the calculations, and it will be clear from context.
For any $Q\in\Delta_{j}$, we denote by $l(Q)=2^{j}$, with the value of $j$ from Theorem 5, and we denote by $r(B)$ the radius of a ball $B$. We denote by $\frac{1}{C}B$ the ball withthe same radius as $B$ but with $1/C$ times the radius of $B$.
@@
We also distinguish between Euclidean cubes with faces parallel to the coordinate planes, such as $\{(x_{1},\dots,x_{n}):b_{i}\leq x_{i}\leq a_{i},\text{for all}\ 1\leq i\leq n.\}$, and the dyadic cubes of David and Semmes[DS1, DS2] mentioned in Theorem 4. We denote by $C(x,r)$ an Euclidean cube with center $x$ and side length $r$.
When $2^{j}\leq C_{2}<2^{j+1}$ for some $j\in\mathbb{Z}$, we write $C_{1}\approx C_{2}$ and $C_{1}=2^{j}$.
@@
$\kappa=\frac{1}{N+1}=\frac{1}{\lceil C_{1}^{2/s}\rceil+1}.$ (16)
@@
Within each such annulus $A_{n}$, consider the gradient at the point $q_{1}$ due to the surface measure within $A_{n}$. We use the condition arising from Eq. 15 above, the Eq. 1 conditions.
Consider the difference in the potential between the points $q_{1},q_{2}\in\partial\Omega$ due to the harmonic measure on the boundary.
@@
Recall that $\mathcal{E}(x,y)$ is the contribution at $x$ due to the fundamental solution, with the pole at $y$. Here we have considered the sets $\mathcal{C}(B(q_{1},t_{1})),\mathcal{C}(B(q_{1},\kappa^{N}t_{1}))$ as the minimal covers of $B(q_{1},t_{1}),B(q_{1},\kappa^{N}t_{1})$ respectively, by dyadic cubes of length $2^{j_{0}}$ for the $j_{0}$ to be determined. Similarly, we define $\mathcal{C}(B(q_{1},t_{1}))$.
@@
Consider the plane perpendicular to the line joining the points $p_{1},q_{1}$, and passing through $q_{1}$ . Call this plane $\Sigma_{1}$. We write $H_{1}^{+}$ for the half plane with the unit outward normal vector $\hat{v_{1}}$ pointing from $q_{1}$ towards $p_{1}$, and the other half plane as $H_{1}^{-}$.
@@
There exists a fixed positive integer $w$ dependent on $\beta$, so that for each $m\geq 0$ there exists a point in $\partial\Omega\cap H^{-}_{1}\cap A_{m}$ which is at a distance at least $c\beta_{1}^{w}\kappa^{m+1}t_{1}$ from $\Sigma_{1}$, and $c$ being some universal constant depending on the dimension $n$. Here, we have,
@@
Consider the disjoint union $\mathcal{D}_{m}$ of $A_{m}\cap\Sigma_{1}$ by Euclidean cubes of radius $c\cdot\kappa^{m+1}t_{1}$. Note from Eq. 16 that we have chosen $\kappa=\frac{1}{\lceil C_{1}^{2/s}\rceil+1}$, and so depending on the parity of $\lceil C_{1}^{2/s}\rceil$, such a $c\geq 1/2$, and such a covering always exists.
@@
To show this, note first that the cube $C(p_{1},r_{1})$ is empty, and the condition Eq. 13, forces the part of $\partial\Omega$ in $H^{+}_{1}\cap A_{m}$ to be much closer to $\Sigma_{1}$ than $\frac{\beta}{2}\kappa^{m}t_{1}$. More precisely, an elementary calculation shows that the maximum distance of $\partial\Omega\cap H^{+}_{1}\cap C(q_{1},\kappa^{m}t_{1})\cap B(p_{1},r_{1})^{c}$ to the plane $\Sigma_{1}$, where $B(p_{1},r_{1})^{c}$ is the complement of the ball $B(p_{1},r_{1})$, is given by $\kappa^{m}t_{1}^{2}/r_{1}$ and we have because of Eq. 13,
@@
Thus, from Eq. 24, we get that there exists some point $y\in C(q_{2},\kappa^{m+1}t_{1})\subset\Sigma_{1}$ so that the cube $C(y,c\kappa^{m+1}t_{1}\beta/2)$ does not contain any point of $\partial\Omega$(1)(1)(1)Here, $c<1$ is a geometric factor dependent on the dimension $n$.. In subsequent iterations of this argument, it might happen that this chosen point $y$ is on the boundary of a cube from the previous generation that is already empty. Even if $y$ is in the interior of the cube $C(q_{2},\kappa^{m+1}t_{1})$, we want to only discard a subcube within $C(y,c\kappa^{m+1}t_{1}\beta/2)\cap C(q_{2},\kappa^{m+1}t_{1})$, of comparable size to $C(y,c\kappa^{m+1}t_{1}\beta/2)$ , that belongs to a covering that is a disjoint union of $C(q_{2},\kappa^{m+1}t_{1})$.
We have some cube $C(q^{\prime\prime},c\kappa^{m+1}t_{1}\beta_{1})\subset C(y,c\kappa^{m+1}t_{1}\beta/2)\cap C(q_{2},\kappa^{m+1}t_{1})$ with $q^{\prime\prime}\in\Sigma_{1}$, that does not contain any point of $\partial\Omega$. Further, because of the choice of $\beta_{1}$, $C(q^{\prime\prime},c\kappa^{m+1}t_{1}\beta_{1})$ is an element of a covering that is a disjoint union of $C(q_{2},\kappa^{m+1}t_{1})$, with radius $c\kappa^{m+1}t_{1}\beta_{1}$. Given the cube $C(q_{2},\kappa^{m+1}t_{1})$, we denote the set of cubes in this covering as $\mathcal{D}_{C(q_{2},\kappa^{m+1}t_{1})}$. We discard all such cubes of the form $C(q^{\prime\prime},c\kappa^{m+1}t_{1}\beta_{1})$, by repeating the same argument over all the cubes of the form $C(q_{2},\kappa^{m+1}t_{1})\in\mathcal{D}_{m}$. We consider the disjoint union of boundary cubes that remains after this process of discarding cubes, within $\Sigma_{1}\cap A_{m}$, and call that $\Sigma_{2}$.
Further, upon a disjoint union over all these cubes in $\mathcal{D}_{m}$ we get a covering of $A_{m}\cap\Sigma_{1}$ by cubes of radius $\kappa^{m+1}t_{1}\beta_{1}$,
@@
and we write $\mathcal{D}^{(1)}_{m}:=\{C(y_{i},c\kappa^{m+1}t_{1}\beta_{1})|y_{i}\in\Sigma_{1}\cap A_{m}\}$.
In particular, the cube $C(q^{\prime\prime},c\kappa^{m+1}t_{1}\beta_{1})$ of the previous paragraph, without loss of generality, is taken to be a cube belonging to $\mathcal{D}_{m}^{(1)}$.
Now, if there is a point in $\partial\Omega\cap H^{-}_{1}\cap A_{m}$ at a distance at least $c\beta\beta_{1}\kappa^{m+1}t_{1}/2\geq c(\beta_{1})^{2}\kappa^{m+1}t_{1}$ away from $\Sigma_{1}$, then also we are done. Otherwise, assume that all points of $\partial\Omega\cap H^{-}_{1}\cap A_{m}$ are at a distance at most $c\kappa^{m+1}t_{1}\beta_{1}\beta/2$ from $\Sigma_{1}$.
By the assumption of the previous paragraph, all points of $\partial\Omega\cap H^{-}_{1}\cap A_{m}$ now belong to the closure of some cube $C(y_{i},c\kappa^{m+1}t_{1}N^{(0)}_{\beta})$ belonging to this collection $\mathcal{D}_{m}^{(1)}$. We now repeat this process. Then for any given cube $C(y_{i},c\kappa^{m+1}t_{1}N^{(0)}_{\beta})\in\mathcal{D}_{m}^{(1)}$ we repeat the previous argument and discard a set of cubes of radius $c^{2}\kappa^{m+1}t_{1}(N^{(0)}_{\beta})^{2}$ . In the process we find a disjoint collection of cubes $\mathcal{D}_{m}^{(2)}$, each now of radius $c^{2}\kappa^{m+1}t_{1}(N^{(0)}_{\beta})^{2}$, each centered on $\Sigma_{1}\cap A_{m}$, and which thus covers the remaining union of cubes that have not been discarded in $\Sigma\cap A_{m}$, and which we define as $\Sigma_{3}$.
In other words, we assume that,
$\sup\limits_{y\in C(x_{i},c\kappa^{m+1}t_{1}N_{\beta}^{(0)})\cap\partial\Omega}\frac{\text{dist}(y,\Sigma_{1})}{c\kappa^{m+1}t_{1}N_{\beta}^{(0)}}\leq\frac{\beta}{2}.$ (28)
@@
$\text{dist}(y,\partial\Omega)\geq\frac{c\kappa^{m+1}t_{1}\beta\beta_{1}}{2}.$ (29)
This gives us, as before, that a cube of the form $C(q_{3},c^{2}\kappa^{m+1}t_{1}(N^{(0)}_{\beta})^{2})$ does not contain any point of $\partial\Omega$, and further that this cube $C(q_{3},c^{2}\kappa^{m+1}t_{1}(N^{(0)}_{\beta})^{2})$ belongs to the disjoint union of cubes that constitute $\Sigma_{3}$.
Now we repeat this argument inductively. Note in particular that the collection of discarded cubes also ensures that the intersection of these cubes with $H^{+}_{1}$ is also empty, because of the fact that the points of $\partial\Omega\cap H^{+}_{1}$ are closer to $\Sigma$ compared to the radius of the cubes under consideration till the stage $w$.
We continue this argument for each positive integer $n$ until we reach a stage where we violate the Ahlfors regularity property of Eq. 2’. Note that we have after the $w$’th stage of the argument, by hypothesis, that all the points of $A_{m}\cap H^{-1}_{1}\cap\partial\Omega$ belong to within a distance at most $\beta\beta_{1}^{w-1}\kappa^{m+1}t_{1}/2$ from $\Sigma_{1}$. In the next step, if there is a point at least a distance $\beta_{1}^{w}\kappa^{m+1}t_{1}/2$ away from $\Sigma_{1}$ within $H^{-1}_{1}$ then we are done, otherwise assume that all these points lie within a distance $\beta_{1}^{w}\kappa^{m+1}t_{1}/2$ from $\Sigma_{1}$. Note that the points in $H^{+}_{1}$ are in any case closer to $\Sigma_{1}$ as quantified by the condition in Eq. 13.
@@
$s\log\kappa+\log((\frac{1}{\kappa})^{n-1}-1)+w(\log(1-(\beta_{1})^{n-1})+(s-n+1)\log\beta_{1})\leq\log(\frac{1}{2}C_{2}^{-2})$ (32)

### v2

@@ Abstract
Given any integer $n\geq 3$, $C_{1}>1$, and $0<\beta<1$, there is a constant $\kappa\in(0,1)$, so that the following holds. Define, $N_{\beta}=\lceil\frac{1}{\beta}\rceil$ and $\beta_{1}:=\frac{1}{4N_{\beta}+2}$.
@@
Note that, $\beta_{1}\leq\beta/4$. In particular, we will only employ the Eq. 1 condition for any point $x\in\mathbb{R}^{n}$, particularly for planes passing through $x$. We do not restrict this condition to only points $x\in\partial\Omega$. Of course, if we were to choose the point $x$ and the plane $V\ni x$ far from the boundary $\partial\Omega$ depending on $\beta$, then the condition is vacuously true, and so the statement is only interesting when the point $x$ and the chosen plane $V$ are close to the boundary.
@@ 3 Preliminaries
$\sup_{y\in C(x,r)\cap\partial\Omega}\frac{\operatorname{dist}(y,V)}{r}\geq\sup_{y\in B(x,r)\cap\partial\Omega}\frac{\operatorname{dist}(y,V)}{r},\sup_{y\in C(x,r)\cap V}\frac{\operatorname{dist}(y,\partial\Omega)}{r}\geq\sup_{y\in B(x,r)\cap V}\frac{\operatorname{dist}(y,\partial\Omega)}{r}$ (10)
@@ 4 Notation
Throughout the proof, $C_{i}>0$ for any $i\geq 1$, denotes a constant whose value may change from line to line. When necessary, we indicate the dependence of $C$ on ambient parameters.
For any $Q\in\Delta_{j}$, we denote by $l(Q)=2^{j}$, with the value of $j$ from Theorem 5, and we denote by $r(B)$ the radius of a ball $B(x,r)$ of radius $r$ centered at $x$. We denote by $\frac{1}{C}B$ the ball withthe same radius as $B$ but with $1/C$ times the radius of $B$. By $S(x,r)$ we denote the sphere of radius $r$ centered at $x$.
@@
We also distinguish between Euclidean cubes with faces parallel to the coordinate planes, such as $\{(x_{1},\dots,x_{n}):b_{i}\leq x_{i}\leq a_{i},\text{for all}\ 1\leq i\leq n.\}$, and the dyadic cubes of David and Semmes[DS1, DS2] mentioned in Theorem 4. As mentioned earlier, $C(x,r)$ is the Euclidean cube given by $C(x,r):=\{(y_{1},\dots,y_{n}):x_{i}-r\leq y_{i}\leq x_{i}+r,1\leq i\leq n\}$, where $x=\{x_{1},\dots,x_{n}\}$. Depending on the context, we will also consider (n1)(n-1)-dimensional cubes contained it hyperplanes.
When $2^{j}\leq C_{2}<2^{j+1}$ for some $j\in\mathbb{Z}$, we write $C_{1}\approx C_{2}$ with $C_{1}=2^{j}$. Throughout this paper, by a covering of any set $E$, we mean a finite set of elements indexed by $J$, $\{E_{j}\}|_{j\geq J}$ so that we have the disjoint union $E=\sqcup_{j\geq 1}E_{j}$ (i.e. for any two $E_{j},E_{k}$ with $j\neq k$, $j,k\in J$ , we have $E_{j}\cap E_{k}=\phi$). Finally, we denote the null set by $\phi$.
@@
$\kappa=\frac{1}{4N+2}=\frac{1}{4\lceil C_{1}^{2/s}\rceil+2}.$ (16)
@@
[…]
Figure 1: The corkscrew point $p_{1}$ and a point $q_{1}\in\partial\Omega$ nearest to $p_{1}$ is shown in the figure. The hyperplane $\Sigma_{1}$ passes through $q_{1}$ and is perpendicular to the line joining $p_{1},q_{1}$, and the cube $C(q_{1},t_{1})$ is also shown. The left and right half planes in the figure are respectively $H^{+}_{1},H^{-}_{1}$. The arc shows a portion of the sphere $S(p_{1},r_{1})$. The dashed lines depict the boundary $\partial\Omega$, which lies in the complement of the ball $B(p_{1},r_{1})$.
@@
Within each such annulus $A_{m}$, consider the gradient at the point $q_{1}$ due to the surface measure within $A_{m}$. We use the condition arising from Eq. 15 above, the Eq. 1 conditions.
Consider the difference in the potential between the points $q_{1},q_{2}\in\partial\Omega$ due to the harmonic measure on $\partial\Omega$.
@@
Recall that $\mathcal{E}(x,y)$ is the contribution at $x$, due to the fundamental solution with the pole at $y$. Here we have considered the sets $\mathcal{C}(B(q_{1},t_{1})),\mathcal{C}(B(q_{1},\kappa^{N}t_{1}))$ as the minimal covers of $B(q_{1},t_{1}),B(q_{1},\kappa^{N}t_{1})$ respectively, by dyadic cubes of length $2^{j_{0}}$ for the $j_{0}$ to be determined(1)(1)(1)See, for example, Eq. 75 for the precise value of $j_{0}$ for the case of s=n1s=n-1, in terms of the $\kappa$ computed in Eq. 16.. Similarly, we define $\mathcal{C}(B(q_{1},t_{1}))$.
@@
Consider the (n1)(n-1)- dimensional plane perpendicular to the line joining the points $p_{1},q_{1}$, and passing through $q_{1}$ . Call this plane $\Sigma_{1}$. We write $H_{1}^{+}$ for the half plane with the unit outward normal vector $\hat{v_{1}}$ pointing from $q_{1}$ towards $p_{1}$, and the other half plane as $H_{1}^{-}$.
@@
There exists a fixed positive integer $w$ dependent on $\beta$, so that for each $m\geq 0$ there exists a point in $\partial\Omega\cap H^{-}_{1}\cap A_{m}$ which is at a distance at least $\beta_{1}^{w}\kappa^{m+1}t_{1}$ from $\Sigma_{1}$.
Here, we have,
@@
For each $m\geq 1$, consider the disjoint union $\mathcal{D}_{m}$ of $A_{m}\cap\Sigma_{1}$ by Euclidean cubes of radius $\kappa^{m+1}t_{1}$. Note from Eq. 16 that we have chosen $\kappa=\frac{1}{4\lceil C_{1}^{2/s}\rceil+2}$, and such a covering always exists. More precisely, we have a disjoint covering $\mathcal{F}_{m}$ of $C(q_{1},\kappa^{m}t_{1})$ by cubes of radius $\kappa^{m+1}t_{1}$ which includes the cube $C(q_{1},\kappa^{m+1}t_{1})$, and we have $\mathcal{D}_{m}=\mathcal{F}_{m}\setminus\{C(q_{1},\kappa^{m+1}t_{1})\}$.
@@
To show this, note first that the ball $B(p_{1},r_{1})$ is empty, and the condition Eq. 13, forces the part of $\partial\Omega$ in $H^{+}_{1}\cap A_{m}$ to be quantifiably closer to $\Sigma_{1}$ than $\frac{\beta}{2}\kappa^{m}t_{1}$. More precisely, an elementary calculation shows that the maximum distance of $\partial\Omega\cap H^{+}_{1}\cap C(q_{1},\kappa^{m}t_{1})\cap B(p_{1},r_{1})^{c}$ to the plane $\Sigma_{1}$, where $B(p_{1},r_{1})^{c}$ is the complement of the ball $B(p_{1},r_{1})$, is given by $\kappa^{m}t_{1}^{2}/r_{1}$ and we have because of Eq. 13,
@@
Thus, from Eq. 24, we get that there exists some point $y\in C(q_{2},\kappa^{m+1}t_{1})\subset\Sigma_{1}$ so that the cube $C(y,\kappa^{m+1}t_{1}\beta/2)$ does not contain any point of $\partial\Omega$(2)(2)(2)Here, $c<1$ is a geometric factor dependent on the dimension $n$.. Here, we mean (n1)(n-1) dimensional cubes contained in $\Sigma_{1}$. Henceforth, this should be clear from context. Note that here and in subsequent iterations of this argument in this lemma, it might happen that this chosen point $y$ is on the boundary of a cube from the previous generation that is already empty. Even if $y$ is in the interior of the cube $C(q_{2},\kappa^{m+1}t_{1})$, we want to only discard a subcube within $C(y,\kappa^{m+1}t_{1}\beta/2)\cap C(q_{2},\kappa^{m+1}t_{1})$, of comparable size to $C(y,\kappa^{m+1}t_{1}\beta/2)$, that belongs to a covering of $C(q_{2},\kappa^{m+1}t_{1})$.
We have some cube $C(q^{\prime\prime},\kappa^{m+1}t_{1}\beta_{1})\subset C(y,\kappa^{m+1}t_{1}\beta/2)\cap C(q_{2},\kappa^{m+1}t_{1})$ with $q^{\prime\prime}\in\Sigma_{1}$, that does not contain any point of $\partial\Omega$. Further, because of the choice of $\beta_{1}$, $C(q^{\prime\prime},\kappa^{m+1}t_{1}\beta_{1})$ is an element of a covering of $C(q_{2},\kappa^{m+1}t_{1})$, with radius $\kappa^{m+1}t_{1}\beta_{1}$. Given the cube $C(q_{2},\kappa^{m+1}t_{1})$, we denote the set of such cubes, of radius $\kappa^{m+1}t_{1}\beta_{1}$, in this covering of $C(q_{2},\kappa^{m+1}t_{1})$, as $\mathcal{D}_{C(q_{2},\kappa^{m+1}t_{1})}$.
Upon a disjoint union over all these cubes in $\mathcal{D}_{m}$ we get a covering of $A_{m}\cap\Sigma_{1}$ by cubes of radius $\kappa^{m+1}t_{1}\beta_{1}$,
@@
We discard all such cubes of the form $C(q^{\prime\prime},\kappa^{m+1}t_{1}\beta_{1})$ chosen in the previous paragraph, by repeating the same argument over all the cubes of the form $C(q_{2},\kappa^{m+1}t_{1})\in\mathcal{D}_{m}$. We consider the disjoint union of boundary cubes that remains after this process, within $\Sigma_{1}\cap A_{m}$, and call that $\Sigma_{2}$. In particular, the cube $C(q^{\prime\prime},\kappa^{m+1}t_{1}\beta_{1})$ belongs to $\mathcal{D}_{m}^{(1)}$, and $C(q^{\prime\prime},\kappa^{m+1}t_{1}\beta_{1})\cap\Sigma_{2}=\phi$ .
Now, if there is a point in $\partial\Omega\cap H^{-}_{1}\cap A_{m}$ at a distance at least $\beta\beta_{1}\kappa^{m+1}t_{1}/2\geq(\beta_{1})^{2}\kappa^{m+1}t_{1}$ away from $\Sigma_{1}$, then also we are done. Otherwise, assume that all points of $\partial\Omega\cap H^{-}_{1}\cap A_{m}$ are at a distance at most $\kappa^{m+1}t_{1}\beta_{1}\beta/2$ from $\Sigma_{1}$.
In other words, we assume that, for any cube $C(y_{i},\kappa^{m+1}t_{1}\beta_{1})\in\mathcal{D}^{(1)}_{m}$, we have
$\sup\limits_{y\in C(y_{i},\kappa^{m+1}t_{1}\beta_{1})\cap\partial\Omega}\frac{\text{dist}(y,\Sigma_{1})}{\kappa^{m+1}t_{1}\beta_{1}}\leq\frac{\beta}{2}.$ (28)
Thus, any point of $\partial\Omega\cap H^{-}_{1}\cap A_{m}$ now belong to the closure of some cube $C(y_{i},\kappa^{m+1}t_{1}\beta_{1})$ belonging to this collection $\mathcal{D}_{m}^{(1)}$. Then for any given cube $C(y_{i},\kappa^{m+1}t_{1}\beta_{1})\in\mathcal{D}_{m}^{(1)}$ we repeat the previous argument and discard a set of cubes of radius $\kappa^{m+1}t_{1}\beta^{2}_{1}$. In the process we find a disjoint collection of cubes $\mathcal{D}_{m}^{(2)}$, each now of radius $c^{2}\kappa^{m+1}t_{1}\beta^{2}_{1}$, each centered on $\Sigma_{1}\cap A_{m}$, and which thus covers the remaining union of cubes that have not been discarded in $\Sigma_{1}\cap A_{m}$, and which we define as $\Sigma_{3}$.
@@
$\text{dist}(y,\partial\Omega)\geq\frac{\kappa^{m+1}t_{1}\beta\beta_{1}}{2}.$ (29)
This gives us, as before, that a cube of the form $C(q_{3},\kappa^{m+1}t_{1}\beta_{1}^{2})$ does not contain any point of $\partial\Omega$, and further that this cube $C(q_{3},\kappa^{m+1}t_{1}\beta_{1}^{2})$ belongs to a covering of $\Sigma_{3}$.
Now we repeat this argument inductively. Note in particular that the argument also ensures that the intersection of these discarded cubes with $H^{+}_{1}$ is also empty, because the points of $\partial\Omega\cap H^{+}_{1}$ are quantifiably closer to $\Sigma$ compared to the radius of the cubes under consideration till the stage $w$, due to Eqs. 25 and 13.
We continue this argument for each positive integer $n$ until we reach a stage where we violate the Ahlfors regularity property of Eq. 2’. Note that we have after the $w$’th stage of the argument, by hypothesis, that all the points of $A_{m}\cap H^{-1}_{1}\cap\partial\Omega$ lie within a distance at most $\beta\beta_{1}^{w-1}\kappa^{m+1}t_{1}/2$ from $\Sigma_{1}$. If there is a point at least a distance $\beta_{1}^{w}\kappa^{m+1}t_{1}/2$ away from $\Sigma_{1}$ within $H^{-1}_{1}$ then we are done, otherwise assume that all these points lie within a distance $\beta_{1}^{w}\kappa^{m+1}t_{1}/2$ from $\Sigma_{1}$. Note that the points in $H^{+}_{1}$ are in any case closer to $\Sigma_{1}$ because of Eqs. 25 and 13.
@@
$s\log\kappa+\log\Big(\big(\frac{1}{\kappa}\big)^{n-1}-1\Big)+w(\log(1-(\beta_{1})^{n-1})+(s-n+1)\log\beta_{1})\leq\log(\frac{1}{2}C_{2}^{-2})$ (32)